The general manager of a chain of furniture stores believes that experience is the most important factor in determining the level of success of a salesperson. To examine this belief she records last month’s sales (in $1,000s) and the years of experience of 10 randomly selected salespeople. These data are listed below.

Salesperson

Years of Experience

Sales ($1,000s)

1

0

7

2

2

9

3

10

20

4

3

15

5

8

18

6

5

14

7

12

20

8

7

17

9

20

30

10

15

25

(a) Given the following data, calculate β1, β0 and determine the regression line (hint #1:calculate β1 first and use y̅ and x̅ in the linear formula to find β0; hint #2: the slope coefficient measures the covariance of x and y accounted for by the variance of x)

n = 10             x̅ = 8.2           y̅ = 17.5       sx2 = 34.76      sy2 = 42.65      sxy = 37.6

(b) Interpret the meaning of the β1 or slope coefficient.

(c) Based on the regression equation and the original raw data, with 10 years of

experience find the predicted level of sales and compare it to the actual level of sales.
Answer #1

scution :- given data 2:10 x =8-2 99.5 Ss = 34.76 sy = 42.65 Szy - 37.6 Required to find requession Equation: Regussion equThe sales will get increased here given by 1.082 c) 8:10 we have f=8.628 +1-082x10 7=19.4487 for actual data , we get for I:1